OpenStudy (anonymous):

If h(x) = x − 1 and j(x) = −4x, find h[j(5)]. −19 −21 −16 −20

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

If f(x) = −x − 1 and g(x) = 4x − 1, find f[g(−2)] −10 −9 −13 8

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

Question 6 (Multiple Choice Worth 1 points) If f(x) = 2x + 1 and g(x) = 3x − 4, find f[g(3)]. 11 17 5 7 Question 7 (Multiple Choice Worth 1 points) If f(x) = −3x − 3 and g(x) = 2x + 1, find (−2). −3 3 −1 1 Question 8 (Multiple Choice Worth 1 points) If f(x) = 2x − 4 and g(x) = x − 2, find f[g(−4)]. 12 −12 −8 8 Question 9 (Multiple Choice Worth 1 points) If r(x) = 3x − 1 and t(x) = −4x − 1, find (r + t)(7). 9 −9 −49 49 Question 10 (Multiple Choice Worth 1 points) If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4). 17 −17 5 −5